# General Riemann Integration on Partially Ordered Net Convergence Structures

### v1

@@ 3.2 Summability in Riesz Spaces and a generalized Kolmogorov Integral
$\sup_{n\in\mathbb{N}}\sum_{i=1}^{n}x_{i}-\sum_{i=1}^{n}x_{i}\leq p_{n},$
$\sum_{j\in J_{0}}x_{j}=\sum_{i=1}^{k}x_{i}\leq\sum_{j\in J}x_{j}.$
Therefore,
Usually we deal with series of non-negative measures valued in Riesz spaces, specially in context with integration, for which we now state a fundamental definition:
Definition 3.13 ($\sigma$-additive Measure in Partially Ordered Spaces).
Let $(\Omega,\mathcal{H})$ a paved space and $(Y,\eta)$ be a Riesz space with Boccuto-Candeloro net convergence structure $\eta$. Then, we say that a non-negative set function:
$\mu:\mathcal{H}\rightarrow Y^{+},$
Its important to notice that the series in the definition above is a unconditional one, because the use of the (commutative operation of) countable union in its definition - a point forcefully put in the topological context by Flesicher and Traynor [64].
By using Propositions 3.5 and 3.6, its possible to prove the following fact directly:
Proposition 3.10 (Monotone Continuity and $\sigma$-Additivity).
Let $(\Omega,\mathcal{H})$ a paved space by a ring and $(Y,\eta)$ be a Riesz space with Boccuto-Candeloro net convergence structure $\eta$. Then, a non-negative finitely additive set function:
$\mu:\mathcal{H}\rightarrow Y^{+},$
$\lim_{(\mathbb{N},\eta)}\mu(A_{n})=0,$
where the limit is decreasing in $n$.
It follows that our definition of $\sigma$-additive measure is the same as the one used in the theory of $\sigma$-additive measures in Riesz spaces equipped with order or (D)-convergence - see [12] for details.
Having a notion of series in partially ordered spaces with abstract convergence structures allow us to define a different type of integral that seems different in nature compared to the Net Riemann integral.
Before the definition itself, consider the following auxiliary object:
Definition 3.14 (Choice Function).
One more convention: for a countable collection of sets $P=\{\alpha\}_{\alpha\in P}$, we denote for $J\in\mathcal{F}(\mathbb{N})$,
$\sum_{j\in J,\alpha\in P}g(\alpha_{j}),$
where $g$ is a set function, to be the sum of the elements of $P$ indexed by $J$.
Now we may define a new integral, which in its original form is due to Kolmogorov [13], but the modern usage of it is due (initially) to Deleanu [14], and Ionescu-Tulcea in [16] and [17], called the $S^{*}$-integral:
Definition 3.15 (Kolmogorov $S^{*}$-integral).
Given and ordered pair:
consisting of
  1. (i)
    $(\Omega,\mathcal{H})$, a paved space $\mathcal{H}$;
  2. (ii)
    $(X,\eta_{1}),(Z,\eta_{2})(Z,\eta_{3})$, Riesz spaces with Boccuto-Candeloro convergence structures connected by a product $\bullet$;
The integral in Definition 3.15 may be paraphrased as follows:
Definition 3.16 ($S^{*}$-integral Unconditional Series).
In other words, whenever the $S^{*}$-integral exists, we assume the existence of the series on the integral for partitions at least cofinal for the convergence (i.e, which refines a fixed element of $\mathcal{P}(\Omega)$).
Now point $(iv)$ of Observation 3.2 may be explained: as the Net Riemann integral uses, in its partially ordered integration system, a finite number of sums for each defining net, the $S^{*}$-integral, defined by a limit of an unconditionally convergent series seems not to be contained in it. We will comment about that later on.
For the rest of the text, we denote by $L_{S^{*}}^{1}(\mu)$ the space of all functions $f:\Omega\rightarrow X$ that are $S^{*}$-integrable with respect to the (fixed) set function $\mu:\Omega\rightarrow Y$ on each $A\in\mathcal{H}$ with the integration structures of Definition 3.15.
We now summarize some properties of the $S^{*}$-integral in this abstract setting, most of them given by Deleanu [14]:
Proposition 3.11 (Properties of the $S^{*}$-Integral).
The $S^{*}$-Integral has the following properties:
  1. (i)
    Let $f,g\in L_{S^{*}}^{1}(\mu)$. Then, for each fixed $A\in\mathcal{H}$
    $\int^{S^{*}}_{A}(f+g)d\mu=\int^{S{*}}_{A}fd\mu+\int^{S{*}}_{A}gd\mu,$
    i.e, the integral is additive on the space $L_{S^{*}}^{1}(\mu)$;
  2. (ii)
    Let $f\in L_{S^{*}}^{1}(\mu)$ such that $f\geq 0$ and $\mu\geq 0$. Then, for each fixed $A\in\mathcal{H}$
    $\int^{S^{*}}_{A}fd\mu\geq 0,$
  3. (iii)
    Let $\mu\geq 0$ and $f,g\in L_{S^{*}}^{1}(\mu)$ such that:
    $g\geq f.$
    Then, for each fixed $A\in\mathcal{H}$
    $\int^{S^{*}}_{A}gd\mu\geq\int^{S^{*}}_{A}fd\mu,$
    i.e, the integral is an isotone map on $L_{S^{*}}^{1}(\mu)$ equipped with the partial ordering inherited from $X$.
  4. (iv)
    Let $\mu\geq 0$ and $f\in L_{S^{*}}^{1}(\mu)$ such that $|f|\in L_{S^{*}}^{1}(\mu)$, then for each fixed $A\in\mathcal{H}$:
    $|\int^{S^{*}}_{A}fd\mu|\leq\int^{S^{*}}_{A}|f|d\mu.$
  5. (v)
    $\int^{S^{*}}_{A\cup B}fd\mu=\int^{S^{*}}_{A}fd\mu+\int^{S^{*}}_{B}fd\mu$
  6. (vi)
  7. (vii)
    $\int_{A}^{S^{*}}fd\mu=c\mu(A)$
For the proof of these results, see [14].
3.3 Examples and Further properties of the (Fleischer) Net Riemann Integral
We now present some examples of the general Net Riemann integral of Definition 3.9. The first three of them can be found, in the topological case and with values in Banach spaces in [21] - its important to notice that all of the examples can be defined in general net convergence monoids, with a partial order or not, but we leave such considerations for a further work.
Notice also that to give examples of the general Net Riemann Integration, and therefore also of the Fleischer Net Riemann integral, is to construct examples of net structures on the (directed) set $\Gamma$ (or $\Gamma_{A}$ for each $A\in\mathcal{H}$ in the case of Fleischer’s integral). For simplicity, we focus on the Net Riemann integral, as the collection of nets for Fleischer Net Convergence integral in Observation 3.1 are obtained directly from the Net Riemann integral case construction. We thus begin with the simplest (abstract) example possible:
@@ 3.3 Examples and Further properties of the (Fleischer) Net Riemann Integral
Let $\mathcal{P}^{f}(\Omega)$ be the set of all finite coverings of a set $\Omega$ in the context of a partially ordered integration structure as in Definition 3.9. Suppose also that the paved space in the definition, $(\Omega,\mathcal{H})$, is a ring paved space such that $\Omega$ is a disjoint union of elements of $\mathcal{H}$  [^6] This is to guarantee that $\Gamma$ in this example is non-empty and directed. We can weaken these and get the same results. This comment is also valid for the next examples.. Then, if we define the refinement partial order $\leq$ on $\mathcal{P}^{f}(\Omega)\times\mathcal{F}^{\mathcal{P}}(\Omega)$ as:
@@
for $P,Q\in\mathcal{P}^{f}(\Omega)$, $m,n\in\mathbb{N}$, and $t_{i}\in P$ , $s_{j}\in Q$, $(\mathcal{P}^{f}(\Omega)\times\mathcal{F}^{\mathcal{P}}(\Omega),\leq)$ becomes a (downwards) directed set (see [21]).
@@
and considering the net formed by the identity, we get that the resulting Net Riemann integral is a refinement net limit (with respect to an abstract Bocutto-Candeloro convergence structrure) of finite sums based on the whole covering $P=\sigma\in\Gamma$. This is a generalization (as we consider coverings, not partitions) of the S-integral of Kolmogorov [13].
@@
Considering different types of refinement type integrals, we may also modify Sion [60] integral to our setting. Fix now the notation as in Definition 3.9.
@@
and considering it equipped with the partial ordering above independent of the chosen points as in Example 3.2 (i.e, the net limit has to be the same for all points or choice functions in the terminology of [21]), it becomes a directed set and the net:
@@
is a Fleischer-Szabó-Száz net defined on it - for more details, such as the proof that $\Gamma$ is a directed set with the partial orderings above, see [21].
In this case, the Net Riemann integral obtained is the integral from [60] which we call the Sion Integral, which in turn constitutes a generalization of the integrals of Rickart [65] and Phillips [66].
We now give an abstract version of the Henstock-Kurzweil (see [67]) integral as in [68].
@@
By the properties $(i)$ and $(ii)$ of the collection $\mathcal{H}$, the directness of $(\Gamma,\leq)$ is a consequence - see [68].
@@
is a Fleischer-Szabó-Száz net which gives the abstract gauge integral of [68].
In general, the relations in $\mathcal{R}$ are given, in specific examples, by gauges (see [67]) in the collection of open sets of the topology of $\Omega$. Thus, the classical gauge integral is contained in this example, as are the integrals of [11], [58], and other integrals of gauge type by specifying the spaces $X,Y,Z$ in Definition 3.9.
We now give an abstract generalization of the construction given in the last example. This is based on the concept of derivation basis in gauge integration, for which our exposition will be based on [69] (see also [7], [6] and the real line case in [70] and [71]). In Henstock [67] and Mcgill [49] terminology, we are going to integrate interval functions on division spaces (therefore, the lattice valued integral of Mcgill [49] is contained in the Net Riemann integral).
@@
We consider now the paved space $(\Omega,\mathcal{H})$ equipped with a measure $\mu$ such that $(\Omega,\mathcal{H},\mu)$ is a measure space - i.e, $\mathcal{H}$ is a $\sigma$-algebra and $\mu$ is a real, positive $\sigma$-additive set function.
We call, following [69], a filterbase $\mathcal{B}$ in the product space $\mathcal{I}\times\Omega$, where $\mathcal{I}$ is a family of sets in $\mathcal{H}$ of positive measure, a derivation basis. We assume now the following conditions on $\mathcal{B}$, which we refer to [69] for details:
@@
In this case, we denote $P_{\beta}(L)$ the set of all $\beta$-partitions of a fixed $\mathcal{B}$-interval $L$. Then, by the partitioning property and the filterbase condition on $\mathcal{B}$, we have that $\{P_{\beta}(L)\}_{\beta\in\mathcal{B}}$ is a filterbase on the product $\mathcal{P}^{f}(L)\times\mathcal{F}(\Omega)$ - [69] again contains all necessary details.
Now, we associate to $\{P_{\beta}(L)\}_{\beta\in\mathcal{B}}$ a Fleischer-Szabó-Száz net. For that, we follow the procedure given in [72] (see also [73] and [74]) for general nets and filterbasis. That is, let
@@
which can be seen to be the same as the ones given in Definition 1 of [69] for the integral (in his notation) $\int_{L}\Phi$. Therefore, his integral is contained in the Net Riemann integral.
@@
We finish this stream of examples with Millington [75] (see also [76]) construction of an abstract integration in terms of runs, an alternative concept to nets and filterbases created by Kenyon and Morse [77] (in fact, with integration theory in mind!). As to not introduce even more concepts in this example, we will translate the results of Millington’s to the language of nets, and also work with a more restricted set of coverings and refer to Millington’s original article for more details concerning the definitions involved here.
@@
For Millington [75] construction, consider $D$ a directed (downwards or upwards) set, $(\Omega,\mathcal{H})$ as in Examples 3.2 and 3.3, $A(\subseteq\Omega)\in\mathcal{H}$ a fixed set, and let $\mathcal{T}$ be the set of all truncations, as in Example 3.3, in $\mathcal{P}(A)$, and $\mathcal{C}$ the set of all choice functions in $\mathcal{H}$ (i.e, $h(C)=c\in C$, for $h:\mathcal{H}\rightarrow\Omega$). Now, consider the net
@@
that is, $(\sigma,\tau)=(\Delta(P_{d}),h_{d,\Delta(P_{d})})$, is a Fleischer-Szabó-Száz net such that the Riemann sums corresponding to it are equal to the second component of the integration run in Definition 2.3 of [75].

### v2

@@ Abstract
The hearth of the result is part $(ii)$ and $(iii)$, which we now prove.
For that, first assume that, given an arbitrary (fixed) set $A\in\mathcal{H}$, $f$ is Pettis integrable on $A$ with (Pettis) integral $x_{A}\in X$. That means that we have, for each $\psi\in\mathcal{L}(X)$,
$\psi(x_{A})=\int^{(P)}_{A}\psi(f)d\mu,$
where the last integral is a Lebesgue integral of the (scalarly measurable) function $\psi(f)$ with respect to $\mu$. Therefore, by hypothesis, $\psi(f)$ is net Riemann integrable with respect to $\mu$ and:
$\int_{A}^{(L)}\psi(f)d\mu=\int_{A}\psi(f)d\mu,$
where the integral in the left side of the equality is a Lebesgue integral, and:
$\int_{A}\psi(f)d\mu=\lim_{(\Gamma(A),\mathbb{R})}S_{\gamma}(\psi(f),\mu),$
where this last limit is taken in the euclidean topology of $\mathbb{R}$, which we indicate by the $\mathbb{R}$ is the subscript.
Now, we prove that the net Riemann integral of $f$ on $A$, in the weak net convergence structure $\eta_{w}$ of $X$, exists and equals $x_{A}$. In this case, we have to prove that, for each $\psi\in\mathcal{L}(X)$, there exists $y_{A}\in X$ such that the following holds:
$y_{A}=\lim_{(\Gamma(A),\eta_{w})}S_{\gamma}(f,\mu),$
that is,
$\psi(y_{A})=\lim_{(\Gamma(A),\mathbb{R})}\psi(S_{\gamma}(f,\mu)),$
for each $\psi\in\mathcal{L}(X)$. But, from the development above, we have that, as $\psi\in\mathcal{L}(X)$ is linear and the sum occurring in the definition of net Riemann integral is finite [^8] Compare this with the integral is Section 3.,
$\psi(S_{\gamma}(f,\mu))=S_{\gamma}(\psi(f),\mu),$
for each $\gamma\in\Gamma(A)$, and the (net) limit of the last sum, in the appropriate convergence structure as above, is exactly $\psi(x_{A})$. Therefore, if we put $y_{A}=x_{A}$, then the net Riemann integral of $f$ with respect to $\mu$ (in the weak net convergence structure of $X$) exists and equals $s_{E}$, where the limit is taken in the weak topology of $X$. Therefore, as $\psi$ and $A$ above were arbitrary, the net Riemann integral in the weak net convergence structure of $X$ exists and is equal to $x_{A}$, for each $A\in\mathcal{H}$. This finishes the first part.
For the converse, assume that, given an arbitrary (fixed) set $A\in\mathcal{H}$, $f$ is net Riemann integrable on $A$ with integral $y_{A}\in X$, where the net limit is taken in the weak net convergence structure of $X$.
Fix $\psi\in\mathcal{L}(X)$ a (continuous) linear functional on $X$. By its continuity, we have that the following limit exists, as consequence of the existence of the net Riemann integral:
$\psi(y_{A})=\lim_{\Gamma(A),\mathbb{R})}\psi(S_{\gamma}(f,\mu)),$
where we identify the right hand side of the expression to be, by definition,
$\int_{A}\psi(f)d\mu,$
which is also, by hypothesis, the Lebesgue integral of $L$ with respect to $\tau$. Therefore, as $\psi$ was arbitrary, if we put $x_{A}=y_{A}$, we have that, for each $\psi\in\mathcal{L}(X)$,
$\psi(y_{A})=\int_{A}\psi(f)d\mu=\int_{A}^{(L)}\psi(f)d\mu,$
which, by uniqueness, proves that $f$ is Pettis integrable on $A$ with Pettis integral equal to the net Riemann integral $y_{A}$. As $A\in\mathcal{H}$ was arbitrary, we conclude that $f$ is Pettis integrable and its Pettis integral equals the net Riemann integral. This finishes the proof, as the other points are immediate. ∎
This shows not only that a general Pettis integration theory can be developed in the net convergence spaces, but that we also may get a Riemmanian representation of such integral. Similar constructions can be made for the Dunford integral in the double dual of net convergence spaces (for the Dunford integral, see Chapter 2 of Sion (1973)), but we are not going to develop it. Some elements of this discussion with respect to the Dunford integral can also be found in Lucyshyn-Wright (2013).
Having now constructed the basic theory of the net Riemann integral, we pass to describe some specific examples of it, and argue that it indeed represents a very interesting generalization of a wide range of Riemann type integrals in the literature (later, we will see the same applied to Lebesgue ones).
2.1 Examples and Further properties of the Net Riemann Integral.
We now present some special cases of the general net Riemann integral of Definition 2.4. The first three of them can be found, in the topological case of Banach spaces in Szabó and Száz (1989). We point out that all of the examples can be defined in general net convergence monoids, with a partial order or not, but we leave such considerations for a further work, and focus in the vector space case.
Notice also that to give examples of the general net Riemann Integration it is only necessary to construct examples of approximating sums in the (directed) sets $\Gamma_{A}$, which should be defined in each procedure for each $A\in\mathcal{H}$. For simplicity, we will generally omit $A$ and work directly with $\Omega$, as the statement for each $A\in\mathcal{H}$ being the same. We thus begin with a simple (abstract) example:
@@
Let $\mathcal{P}^{f}(\Omega)$ be the set of all finite coverings of a set $\Omega$ in the context of an integration structure as in Definition 2.4. Suppose also that the paved space in the definition, $(\Omega,\mathcal{H})$, is a ring paved space such that $\Omega$ is a disjoint union of elements of $\mathcal{H}$  [^9] This is to guarantee that $\Gamma$ in this example is non-empty and directed. We can weaken these and get the same results. This comment is also valid for the next examples.. Then, if we define the refinement partial order $\leq$ on $\mathcal{P}^{f}(\Omega)\times\mathcal{F}^{\mathcal{P}}(\Omega)$ as:
@@
for $P,Q\in\mathcal{P}^{f}(\Omega)$, $m,n\in\mathbb{N}$, and $t_{i}\in P$ , $s_{j}\in Q$, $(\mathcal{P}^{f}(\Omega)\times\mathcal{F}^{\mathcal{P}}(\Omega),\leq)$ becomes a (downwards) directed set (see Szabó and Száz (1989)).
@@
and considering the net formed by the identity, we get that the resulting Net Riemann integral is a refinement net limit (with respect to the net convergence structure of $Z$) of finite sums based on the whole covering $P=\sigma\in\Gamma$. This is a generalization (as we consider coverings, not partitions) of the S-integral of Kolmogorov Kolmogoroff (1930).
@@
Considering different types of refinement type integrals, we may also modify Sion (1973) integral to our setting. Fix now the notation as in Definition 2.4.
@@
and considering it equipped with the partial ordering above independent of the chosen points as in Example 2.1 (i.e, the net limit has to be the same for all points or choice functions in the terminology of Szabó and Száz (1989)), it becomes a directed set and the net:
@@
is a Fleischer-Szabó-Száz net defined on it - for more details, such as the proof that $\Gamma$ is a directed set with the partial orderings above, see Szabó and Száz (1989).
In this case, the Net Riemann integral obtained is the integral from Sion (1973) which we call the Sion Integral, which in turn constitutes a generalization of the classical Rickart and Phillips integrals of Rickart (1942) and Phillips (1940).
We now give an abstract version of the Henstock-Kurzweil (see Henstock (1991)) integral as in Száz (1993).
@@
By the properties $(i)$ and $(ii)$ of the collection $\mathcal{H}$, the directness of $(\Gamma,\leq)$ is a consequence - see Száz (1993).
@@
is a Fleischer-Szabó-Száz net which gives the abstract gauge integral of Száz (1993).
In general, the relations in $\mathcal{R}$ are given, in specific examples, by gauges (see Henstock (1991)) in the collection of open sets of the topology of $\Omega$. Thus, the classical gauge integral is contained in this example, as are the integrals of Boccuto et al. (2012b), Riečan and Vrábelová (1993), and other integrals of gauge type by specifying the spaces $X,Y,Z$ in Definition 2.4.
We now give an abstract generalization of the construction given in the last example. This is based on the concept of derivation basis in gauge integration, for which our exposition will be based on Skvortsov (2025) (see also Boccuto and Skvortsov (2006), Boccuto and Skvortsov (2004) and the real line case in Thomson (1983a) and Thomson (1983b)). In Henstock’s and Mcgill’s terminologies of Henstock (1991) and McGill (1975) respectively, we are going to integrate interval functions on division spaces (therefore, the lattice valued integral of McGill (1975) is contained in the Net Riemann integral).
@@
We consider now the paved space $(\Omega,\mathcal{H})$ equipped with a measure $\mu$ such that $(\Omega,\mathcal{H},\mu)$ is a measure space - i.e, $\mathcal{H}$ is a $\sigma$-algebra and $\mu$ is a real, positive $\sigma$-additive set function  [^10] We can actually weaken these hypothesis and consider the general group valued integral in Chapter 2 of Henstock (1991). Nevertheless, the procedure is the same, so we omit further details..
We call, following Skvortsov (2025), a filterbase $\mathcal{B}$ in the product space $\mathcal{I}\times\Omega$, where $\mathcal{I}$ is a family of sets in $\mathcal{H}$ of positive measure, a derivation basis. We assume now the following conditions on $\mathcal{B}$, which we refer to Skvortsov (2025) for details:
@@
In this case, we denote $P_{\beta}(L)$ the set of all $\beta$-partitions of a fixed $\mathcal{B}$-interval $L$. Then, by the partitioning property and the filterbase condition on $\mathcal{B}$, we have that $\{P_{\beta}(L)\}_{\beta\in\mathcal{B}}$ is a filterbase on the product $\mathcal{P}^{f}(L)\times\mathcal{F}(\Omega)$ - Skvortsov (2025) again contains all necessary details.
Now, we associate to $\{P_{\beta}(L)\}_{\beta\in\mathcal{B}}$ a Fleischer-Szabó-Száz net. For that, we follow the procedure given in Costantini and Priola (2001) (see also Bartle (1963) and Kennedy (1984)) for general nets and filterbasis. That is, let
@@
which can be seen to be the same as the ones given in Definition 1 of Skvortsov (2025) for the integral (in his notation) $\int_{L}\Phi$. Therefore, his integral is contained in the Net Riemann integral.
@@
We finish this stream of examples with Millington integral of Millington (2005) (see also Millington (2007)), constructed as an abstract integration in terms of runs, an alternative concept to nets and filterbasis created by Kenyon and Morse (1958) (in fact, with integration theory in mind!). As to not introduce even more concepts in this example, we will translate the results of Millington’s to the language of nets, and also work with a more restricted set of coverings and refer to Millington’s original article for more details concerning the definitions involved here.
@@
For Millington (2005) construction, consider $D$ a directed (downwards or upwards) set, $(\Omega,\mathcal{H})$ as in Examples 2.1 and 2.2, $A(\subseteq\Omega)\in\mathcal{H}$ a fixed set, and let $\mathcal{T}$ be the set of all truncations, as in Example 2.2, in $\mathcal{P}(A)$, and $\mathcal{C}$ the set of all choice functions in $\mathcal{H}$ (i.e, $h(C)=c\in C$, for $h:\mathcal{H}\rightarrow\Omega$). Now, consider the net
@@
that is, $(\sigma,\tau)=(\Delta(P_{d}),h_{d,\Delta(P_{d})})$, is a Fleischer-Szabó-Száz net such that the Riemann sums corresponding to it are equal to the second component of the integration run in Definition 2.3 of Millington (2005).